OpenStudy (anonymous):

What is the common ratio of this geometric sequence? 1 27 2 18 3 12 4 8 5 16/3 A. -9 B. 2/3 C. 3/2 D. 9

OpenStudy (the_fizicx99):

Divide the succeeding term by the previous term, 18 / 27 12 / 18 8 / 12

OpenStudy (anonymous):

Test taking strategies-- A. -9: definitely not because it isn't changing that much B. 2/3: possibly, check it in a second C. 3/2: If the common ratio is greater than 1, then the terms in the sequence would be getting bigger--but they're getting smaller, so it can't be C D. 9: no, for the same reason A doesn't work. None of those terms is nine times larger than any other in magnitude. So B has to be the answer, but just to check: 27*2/3 = 18 18*2/3 = 12 12*2/3 = 8 8*2/3 = 16/3 So B must be the answer